# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ 1.3 Filtering
$=\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right].$ (100)
The second derivative of $\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ with respect to $\boldsymbol{\theta}^{i}_{t}$ is given by:
$\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})}{\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right)$ $=\sum_{l=1}^{L}\left[-\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right)\right]-(\mathbf{W}^{i}_{t|t-1})^{-1}$ (101)
@@
$\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}(\boldsymbol{\theta}^{i}_{t})^{\top}}$ $=\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right]\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}$
$=\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right]$
$\phantom{=}\times\left[\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}}\right]\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}$
$=r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\{1-r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}$ (102)
Using the first and second-order derivatives, the MAP estimate $\boldsymbol{\theta}^{i}_{\rm MAP}$ was found by the Newton-Raphson method.
After finding the MAP estimate, we approximate the filter distribution by a Gaussian distribution,
$p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}^{1:L}_{1:t},\mathbf{w})\simeq N(\boldsymbol{\theta}^{i}_{t|t},\mathbf{W}^{i}_{t|t})\propto\exp[(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{\rm MAP})\mathbf{W}^{-1}_{t|t}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{\rm MAP})]$ (103)
with the following mean and variance via the Laplace approximation:
$\boldsymbol{\theta}^{i}_{t|t}=\boldsymbol{\theta}_{\rm MAP},$ (104)
@@
$\mathbf{W}^{i}_{t|t}$ $=\left[-\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial\log p\left(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{1-t},\mathbf{w}\right)}{\partial\left(\boldsymbol{\theta}^{i}_{t}\right)^{\mathsf{T}}}\right)\bigg|_{\boldsymbol{\theta}^{i}_{t}=\boldsymbol{\theta}_{\rm MAP}}\right]^{-1}$
$=\left[\mathbf{G}\left(\boldsymbol{\theta}_{\rm MAP}\right)+\left(\mathbf{W}^{i}_{t|t-1}\right)^{-1}\right]^{-1},$ (105)
where $\mathbf{G}^{i}(\boldsymbol{\theta})$ is given by
$\mathbf{G}\left(\boldsymbol{\theta}^{i}_{t}\right)$ $=\sum_{l=1}^{L}\left[\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right)\right]$
$=\sum_{l=1}^{L}r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\{1-r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}.$ (106)
By sequentially applying the one-step prediction density and the filter density sequentially for $t=1,\ldots,T$, we obtain the filter densities of all time steps.
@@ 1.4 Smoothing
Given that the filter density is approximated by Gaussian distributions, the smoothing distribution for the parameters of each neuron can be computed iteratively by using the filter density and the one-step prediction density in a backward pass from the final time step $T$, following the Rauch-Tung-Striebel smoother [96]:
$\mathbf{W}^{i}_{t-1|T}$ $=\mathbf{W}^{i}_{t-1|t-1}+\mathbf{A}^{i}_{t-1}\left(\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t|t-1}\right)\left(\mathbf{A}^{i}_{t-1}\right)^{\top},$ (108)
$\mathbf{A}^{i}_{t-1}$ $=\mathbf{W}^{i}_{t-1|t-1}\left(\mathbf{W}^{i}_{t|t-1}\right)^{-1},\quad t=2,\ldots,T.$ (109)
For completeness, we provide a compact derivation of these equations below.
@@
$=\int p(\boldsymbol{\theta}^{i}_{t-1}|\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{0:t-1},\mathbf{w})p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:T},\mathbf{w})\,d\boldsymbol{\theta}^{i}_{t}.$ (110)
Here we used the Markovian assumption at the second equality. The condtional density $p(\boldsymbol{\theta}^{i}_{t-1}|\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{0:t-1},\mathbf{w})$ is obtained as
@@
$=\frac{p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1}\mathbf{w})p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1},\mathbf{w})}{p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})},$ (111)
which is composed of the filter and one-step prediction densities and the state model. Since we asssume that these are Gaussian distributions, given that the smoother density at time $t$ is Gaussian, the linear operations in Eqs. 110 and 111 gurantees that the smoother density at time t1t-1 is Gaussian. Therefore, the distribution is specified by the mean and covarinace defined as
$\boldsymbol{\theta}^{i}_{t-1|T}$ $\equiv E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:T}}\left[\boldsymbol{\theta}^{i}_{t-1}\right]$ (112)
$\mathbf{W}^{i}_{t-1|T},$ $\equiv E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:T}}\left[\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|T}\right)\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|T}\right)^{\top}\right].$ (113)
To obtain their closed form expressions, first we note that the joint denisty in Eq. 111 is written as
$p(\boldsymbol{\theta}^{i}_{t-1},\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})=\mathcal{N}\left(\begin{pmatrix}\boldsymbol{\theta}^{i}_{t-1|t-1}\\ \boldsymbol{\theta}^{i}_{t|t-1}\end{pmatrix},\begin{pmatrix}\mathbf{W}^{i}_{t-1|t-1}&\mathbf{W}^{i}_{t-1,t|t-1}\\ \mathbf{W}^{i}_{t,t-1|t-1}&\mathbf{W}^{i}_{t|t-1}\end{pmatrix}\right)$, (114)
where $\mathbf{W}^{i}_{t-1,t|t}$ is the cross covariance given the data up to time t1t-1. Here, we note that, under the linear Gaussian transition with an identity transition matrix, the one-step prediciton mean is
$\boldsymbol{\theta}^{i}_{t|t-1}=\boldsymbol{\theta}^{i}_{t-1|t-1}$ (115)
@@
$\mathbf{W}^{i}_{t-1,t|t}$ $\equiv E_{\boldsymbol{\theta}^{i}_{t-1},\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t}}\left[\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|t-1}\right)\left(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1}\right)^{\top}\right]$
$=E_{\boldsymbol{\theta}^{i}_{t-1},\boldsymbol{\xi}_{t}|\mathbf{x}_{0:t}}\left[\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|t-1}\right)\left(\boldsymbol{\theta}^{i}_{t-1}+\boldsymbol{\xi}_{t}-\boldsymbol{\theta}^{i}_{t-1|t-1}\right)^{\top}\right]$
@@
Given the joint density, we obtain the condtional density (Eq. 111). We note that given the multivariate normal distirbution,
@@
The conditional distribution of $\mathbf{x}_{a}\mid\mathbf{x}_{b}$ is
$\mathbf{x}_{a}\mid\mathbf{x}_{b}\sim\mathcal{N}(\boldsymbol{\mu}_{a|b},\boldsymbol{\Sigma}_{a|b})$ (118)
@@
$\boldsymbol{\mu}_{a|b}$ $=\boldsymbol{\mu}_{a}+\boldsymbol{\Sigma}_{ab}\boldsymbol{\Sigma}_{bb}^{-1}(\mathbf{x}_{b}-\boldsymbol{\mu}_{b})$ (119)
$\boldsymbol{\Sigma}_{a|b}$ $=\boldsymbol{\Sigma}_{aa}-\boldsymbol{\Sigma}_{ab}\boldsymbol{\Sigma}_{bb}^{-1}\boldsymbol{\Sigma}_{ba}.$ (120)
@@
$p(\boldsymbol{\theta}^{i}_{t-1}|\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{0:t-1},\mathbf{w})$ $=\mathcal{N}\left(\boldsymbol{\theta}^{i}_{t-1|t-1}+\mathbf{A}_{t-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1|t-1}),\right$.
$\phantom{====}\left.\mathbf{W}^{i}_{t-1|t-1}-\mathbf{A}_{t-1}\mathbf{W}^{i}_{t-1|t-1}\right),$ (121)
@@
Finally, the smoothing density at time $t$ is obtained by multiplying the smoother density at time $t$ and integrating out $\boldsymbol{\theta}^{i}_{t}$ according to Eq. 110. For this, we note that, given the following two normal distributsions:
$p(\mathbf{x}_{a}\mid\mathbf{x}_{b})$ $=\mathcal{N}(\mathbf{x}_{a}\mid\mathbf{A}\mathbf{x}_{b}+\mathbf{b},\boldsymbol{\Sigma}_{a|b}),$ (122)
$p(\mathbf{x}_{b})$ $=\mathcal{N}(\mathbf{x}_{b}\mid\boldsymbol{\mu}_{b},\mathbf{\Sigma}_{b}),$ (123)
@@
$p(\mathbf{x}_{a})=\int p(\mathbf{x}_{a}\mid\mathbf{x}_{b})\,p(\mathbf{x}_{b})\,d\mathbf{x}_{b}=\mathcal{N}(\mathbf{x}_{a}\mid\boldsymbol{\mu}_{a},\boldsymbol{\Sigma}_{a}),$ (124)
@@
$\boldsymbol{\mu}_{a}=\mathbf{A}\boldsymbol{\mu}_{b}+\mathbf{b},\qquad\boldsymbol{\Sigma}_{a}=\mathbf{A}\mathbf{\Sigma}_{b}\mathbf{A}^{\top}+\boldsymbol{\Sigma}_{a|b}.$ (125)
Applying this formula to Eq. 121 and $p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:T},\mathbf{w})=\mathcal{N}\left(\boldsymbol{\theta}^{i}_{t|T},\mathbf{W}^{i}_{t|T}\right)$, we obtain the smoothing density $p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:T},\mathbf{w})$ whose mean and covariance are given by
$\boldsymbol{\theta}^{i}_{t-1|T}$ $=\mathbf{A}_{t-1}\boldsymbol{\theta}^{i}_{t|T}+\left(\boldsymbol{\theta}^{i}_{t-1|t-1}-\mathbf{A}_{t-1}\boldsymbol{\theta}^{i}_{t-1|t-1}\right)$
$=\boldsymbol{\theta}^{i}_{t-1|t-1}+\mathbf{A}_{t-1}(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|t-1}),$ (126)
@@
$=\mathbf{W}^{i}_{t-1|t-1}+\mathbf{A}_{t-1}\left(\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t|t-1}\right)\mathbf{A}_{t-1}^{\top}.$ (127)
@@
1.5 Optimization of parameters
We consider the problem for optimizing the parameter that maximizes the marginal likelihood function. Instead of the computational expensive marginal likelihood, we optimize its lower bound using Expectation-Maximization (EM) algorithm. In the EM algorithm, at E-step, the posterior density is obtained under given hyperparameters via the algorithm described in the previous section. Then, we optimize the hyperparameters that maximizes the lower bound, using the given posterior density at M-step. Using Jensen’s inequality $\log E[X]\geq E[\log X]$, this lower bound is given by

### v3

@@ 1.3 Filtering
$=\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right].$ (S1.24)
The second derivative of $\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ with respect to $\boldsymbol{\theta}^{i}_{t}$ is given by
$\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})}{\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right)$ $=\sum_{l=1}^{L}\left[-\frac{\partial^{2}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}\right]-(\mathbf{W}^{i}_{t|t-1})^{-1}.$ (S1.25)
@@
$\frac{\partial^{2}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}(\boldsymbol{\theta}^{i}_{t})^{\top}}$ $=\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right]\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}$
$=\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right]\left[\mathbf{F}(1,\mathbf{x}^{l}_{t-1})-\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}}\right]\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}$
$=r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\{1-r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}.$ (S1.26)
Using the first and second-order derivatives, the MAP estimate $\boldsymbol{\theta}^{i}_{\rm MAP}$ for each neurons was found by the Newton-Raphson method.
After finding the MAP estimate, we approximate the filter density by a Gaussian distribution via the Laplace’s method,
$p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{1:t},\mathbf{w})=\frac{1}{\sqrt{|2\pi\mathbf{W}_{t|t}|}}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})^{\top}\mathbf{W}_{t|t}^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t})\right]$ (S1.27)
with the following mean and variance:
$\boldsymbol{\theta}^{i}_{t|t}=\boldsymbol{\theta}^{i}_{\rm MAP},$ (S1.28)
@@
$\mathbf{W}^{i}_{t|t}$ $=\left[-\frac{\partial}{\partial\boldsymbol{\theta}^{i}_{t}}\left(\frac{\partial\log p\left(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{1-t},\mathbf{w}\right)}{\partial\left(\boldsymbol{\theta}^{i}_{t}\right)^{\mathsf{T}}}\right)\bigg|_{\boldsymbol{\theta}^{i}_{t}=\boldsymbol{\theta}^{i}_{t|t}}\right]^{-1}$
$=\left[\mathbf{G}\left(\boldsymbol{\theta}^{i}_{t|t}\right)+\left(\mathbf{W}^{i}_{t|t-1}\right)^{-1}\right]^{-1},$ (S1.29)
where $\mathbf{G}\left(\boldsymbol{\theta}^{i}_{t}\right)$ is given by
$\mathbf{G}\left(\boldsymbol{\theta}^{i}_{t}\right)$ $=\sum_{l=1}^{L}\frac{\partial^{2}\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}\partial(\boldsymbol{\theta}^{i}_{t})^{\top}}$
$=\sum_{l=1}^{L}r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\{1-r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})\mathbf{F}(1,\mathbf{x}^{l}_{t-1})^{\top}.$ (S1.30)
By sequentially applying the one-step prediction density and the filter density for $t=1,\ldots,T$, we obtain the filter densities of all time steps.
@@ 1.4 Smoothing
Given that the filter density is approximated by Gaussian distributions, the smoothing density for the parameters of each neuron can be computed iteratively by using the filter density and the one-step prediction density in a backward manner from the final time step $T$, following the Rauch-Tung-Striebel smoother [1]:
$\boldsymbol{\theta}^{i}_{t-1|T}$ $=\boldsymbol{\theta}^{i}_{t-1|t-1}+\mathbf{A}^{i}_{t-1}\left(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t|t}\right),$ (S1.31)
$\mathbf{W}^{i}_{t-1|T}$ $=\mathbf{W}^{i}_{t-1|t-1}+\mathbf{A}^{i}_{t-1}\left(\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t|t-1}\right)\left(\mathbf{A}^{i}_{t-1}\right)^{\top},$ (S1.32)
$\mathbf{A}^{i}_{t-1}$ $=\mathbf{W}^{i}_{t-1|t-1}\left(\mathbf{W}^{i}_{t|t-1}\right)^{-1},$ (S1.33)
for $t=2,\ldots,T$. For completeness, we provide a compact derivation of these equations below.
@@
$=\int p(\boldsymbol{\theta}^{i}_{t-1}|\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{0:t-1},\mathbf{w})p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:T},\mathbf{w})\,d\boldsymbol{\theta}^{i}_{t}.$ (S1.34)
Here, we used the Markovian assumption at the second equality. The conditional density $p(\boldsymbol{\theta}^{i}_{t-1}|\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{0:t-1},\mathbf{w})$ is obtained as
@@
$=\frac{p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{w})p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1},\mathbf{w})}{p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})},$ (S1.35)
which is composed of the filter and one-step prediction densities, and the state model. Since we assume that these are Gaussian distributions, given that the smoother density at time $t$ is Gaussian, the linear operations in Eqs. S1.34 and S1.35 guarantee that the smoother density at time t1t-1 is Gaussian. Therefore, the distribution is specified by the mean and covariance defined as
$\boldsymbol{\theta}^{i}_{t-1|T}$ $\equiv E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:T}}\boldsymbol{\theta}^{i}_{t-1}$ (S1.36)
$\mathbf{W}^{i}_{t-1|T}$ $\equiv E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:T}}\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|T}\right)\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|T}\right)^{\top}.$ (S1.37)
To obtain their closed form expressions, first we note that the joint density in Eq. S1.35 is written as
$p(\boldsymbol{\theta}^{i}_{t-1},\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t-1},\mathbf{w})=\mathcal{N}\left(\begin{pmatrix}\boldsymbol{\theta}^{i}_{t-1}\\ \boldsymbol{\theta}^{i}_{t}\end{pmatrix};\begin{pmatrix}\boldsymbol{\theta}^{i}_{t-1|t-1}\\ \boldsymbol{\theta}^{i}_{t|t-1}\end{pmatrix},\begin{pmatrix}\mathbf{W}^{i}_{t-1|t-1}&\mathbf{W}^{i}_{t-1,t|t-1}\\ \mathbf{W}^{i}_{t,t-1|t-1}&\mathbf{W}^{i}_{t|t-1}\end{pmatrix}\right)$, (S1.38)
where $\mathbf{W}^{i}_{t-1,t|t-1}$ is the cross covariance given the data up to time t1t-1. Here, we note that, under the linear Gaussian transition with an identity transition matrix, the one-step prediction mean is
$\boldsymbol{\theta}^{i}_{t|t-1}=\boldsymbol{\theta}^{i}_{t-1|t-1}.$ (S1.39)
@@
$\mathbf{W}^{i}_{t-1,t|t}$ $\equiv E_{\boldsymbol{\theta}^{i}_{t-1},\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:t}}\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|t-1}\right)\left(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1}\right)^{\top}$
$=E_{\boldsymbol{\theta}^{i}_{t-1},\boldsymbol{\xi}_{t}|\mathbf{x}_{0:t}}\left(\boldsymbol{\theta}^{i}_{t-1}-\boldsymbol{\theta}^{i}_{t-1|t-1}\right)\left(\boldsymbol{\theta}^{i}_{t-1}+\boldsymbol{\xi}_{t}-\boldsymbol{\theta}^{i}_{t-1|t-1}\right)^{\top}$
@@
Given the joint density, we obtain the conditional density (Eq. S1.35). We note that given the multivariate normal distribution,
@@
The conditional distribution of $\mathbf{x}_{a}{|}\mathbf{x}_{b}$ follows
$\mathbf{x}_{a}{|}\mathbf{x}_{b}\sim\mathcal{N}(\boldsymbol{\mu}_{a|b},\boldsymbol{\Sigma}_{a|b})$ (S1.42)
@@
$\boldsymbol{\mu}_{a|b}$ $=\boldsymbol{\mu}_{a}+\boldsymbol{\Sigma}_{ab}\boldsymbol{\Sigma}_{bb}^{-1}(\mathbf{x}_{b}-\boldsymbol{\mu}_{b}),$ (S1.43)
$\boldsymbol{\Sigma}_{a|b}$ $=\boldsymbol{\Sigma}_{aa}-\boldsymbol{\Sigma}_{ab}\boldsymbol{\Sigma}_{bb}^{-1}\boldsymbol{\Sigma}_{ba}.$ (S1.44)
@@
$p(\boldsymbol{\theta}^{i}_{t-1}|\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{0:t-1},\mathbf{w})$ $=\mathcal{N}\left(\boldsymbol{\theta}^{i}_{t-1};\boldsymbol{\theta}^{i}_{t-1|t-1}+\mathbf{A}_{t-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1|t-1}),\mathbf{W}^{i}_{t-1|t-1}-\mathbf{A}_{t-1}\mathbf{W}^{i}_{t-1|t-1}\right),$ (S1.45)
@@
Finally, the smoothing density at time $t$ is obtained by multiplying the smoother density at time $t$ and integrating out $\boldsymbol{\theta}^{i}_{t}$ according to Eq. S1.34. For this, we note that, given the following two normal distributions:
$p(\mathbf{x}_{a}{|}\mathbf{x}_{b})$ $=\mathcal{N}(\mathbf{x}_{a};\mathbf{A}\mathbf{x}_{b}+\mathbf{b},\boldsymbol{\Sigma}_{a|b}),$ (S1.46)
$p(\mathbf{x}_{b})$ $=\mathcal{N}(\mathbf{x}_{b};\boldsymbol{\mu}_{b},\boldsymbol{\Sigma}_{b}),$ (S1.47)
@@
$p(\mathbf{x}_{a})=\int p(\mathbf{x}_{a}{|}\mathbf{x}_{b})\,p(\mathbf{x}_{b})\,d\mathbf{x}_{b}=\mathcal{N}(\mathbf{x}_{a};\boldsymbol{\mu}_{a},\boldsymbol{\Sigma}_{a}),$ (S1.48)
@@
$\boldsymbol{\mu}_{a}$ $=\mathbf{A}\boldsymbol{\mu}_{b}+\mathbf{b},$ (S1.49)
$\boldsymbol{\Sigma}_{a}$ $=\mathbf{A}\boldsymbol{\Sigma}_{b}\mathbf{A}^{\top}+\boldsymbol{\Sigma}_{a|b}.$ (S1.50)
Applying this formula to Eq. S1.45 and $p(\boldsymbol{\theta}^{i}_{t}|\mathbf{x}_{0:T},\mathbf{w})=\mathcal{N}\left(\boldsymbol{\theta}^{i}_{t};\boldsymbol{\theta}^{i}_{t|T},\mathbf{W}^{i}_{t|T}\right)$, we obtain the smoothing density $p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:T},\mathbf{w})$ whose mean and covariance are given by
$\boldsymbol{\theta}^{i}_{t-1|T}$ $=\boldsymbol{\theta}^{i}_{t-1|t-1}+\mathbf{A}_{t-1}(\boldsymbol{\theta}^{i}_{t|T}-\boldsymbol{\theta}^{i}_{t-1|t-1}),$ (S1.51)
@@
$=\mathbf{W}^{i}_{t-1|t-1}+\mathbf{A}_{t-1}\left(\mathbf{W}^{i}_{t|T}-\mathbf{W}^{i}_{t|t-1}\right)\mathbf{A}_{t-1}^{\top}.$ (S1.52)
@@
1.5 Optimization of hyperparameters
We consider the problem of optimizing the hyperparameters that maximize the marginal likelihood function. Instead of the marginal likelihood, we optimize its tractable lower bound. In the Expectation-Maximization (EM) algorithm, the posterior density is obtained under given hyperparameters via the algorithm described in the previous section at the E-step. At the M-step, we optimize the hyperparameters that maximize the lower bound, using the given posterior density. Using Jensen’s inequality $\log E[X]\geq E[\log X]$, this lower bound is given by